# Tensor-Var: Variational Data Assimilation in Tensor Product Feature Space

### v2

@@ Appendix A Theoretical Analysis
In this section, we provide a theoretical convergence analysis of Tensor-Var, drawing on concepts from control theory and contraction analysis. We begin by introducing comparison functions, including class $\mathcal{K}$ and $\mathcal{KL}$ functions, as well as Lyapunov functions. Using these tools, we examine the convergence of Tensor-Var through a differential equation, demonstrating monotonic contraction based on Lyapunov direct method. Furthermore, by using comparison functions, we show that contraction in the feature space implies contraction in the original space.
@@
Assumption 1. To conduct a formal convergence analysis of Tensor-Var, we make a mild assumption regarding the first-order differentiability of the dynamical system $F$ in equation 1 with time derivative $\dot{s}=f(s)$, a standard assumption in convergence studies (Sastry, 2013).
@@
A.1 Notations and Technical Lemmas
Definition A.1 (Class $\mathcal{K}$ function (Gajic & Qureshi, 2008).).
@@ A.1 Notations and Technical Lemmas
Definition A.2 (Class $\mathcal{KL}$ function (Gajic & Qureshi, 2008).).
A continuous function $\beta:[0,a]\times[0,\infty)\to[0,\infty)$ is said to belong to class $\mathcal{KL}$ if each fixed $t$, the mapping $\beta(r,t)$ belonging to the class $\mathcal{K}$ with respect to $r$ and, for each fixed $a$ the mapping $\beta(a,t)$ is decreasing with respect to $t$ and, $\beta(a,t)\to 0$ as $t\to\infty$.
The $\mathcal{K}$ and $\mathcal{KL}$ are two classes of comparison functions, we can use the comparison function to analyze the monotone contraction in both spatial and temporal horizons.
Definition A.3 (Lyapunov Stability (Gajic & Qureshi, 2008)).
If the Lyapunov function $V$ is globally positive definite, radially unbounded, the equilibrium isolated and the time derivative of the Lyapunov function is globally negative definite:
@@
then the equilibrium is proven to be globally asymptotically stable. The Lyapunov function is a class $\mathcal{K}$ function, which satisfying the condition as follow
@@
Lemma A.1 (Hurwitz stability criterion (Parks, 1962)).
@@
Lemma A.2 (Hurwitz stability criterion via Lyapunov function (Sastry, 2013)).
@@
Definition A.4 (Contraction Manchester & Slotine (2017)).
@@
A.2 Proof of Key Theorems
To adopt the convergence analysis of DA problem, we give a mild assumption on the smoothness of dynamics, the first-order derivative exists in the system 1.
Theorem A.3 (Embedding and Consistent convergence).
@@ A.2 Proof of Key Theorems
The second line of equation LABEL:Equation:_dynamical_system_in_Hilbert_space follows from chain rule, the final line represents the time derivative in Hilbert space, where $\mathcal{A}$ is a linear operator that governs the dynamics. Following the work from Romanoff (1947); Bobrowski (2016), we define $\mathcal{A}$ as:
@@
where $\mathcal{C}_{S^{+}\mid S}$ is the conditional covariance operator between future and current states in the RKHS. Given the smoothness of the kernel function and the differentiability of the system dynamics, the linear operator $\mathcal{A}$ exists and well-defined in this context.
Since $s_{*}$ is the equilibrium point, we can derive a natural result that $f(s_{*})=0$. Invoking equation LABEL:Equation:_dynamical_system_in_Hilbert_space (third line) we have $\frac{d\phi_{S}}{dt}(s_{*})=0$. Thus $\phi_{S}(s_{*})$ is also a local equilibrium point in RKHS. From the embedding property of RKHS, we have a local injective map $\phi:\mathbb{R}^{n_{s}}\to\mathbb{R}^{d_{s}}$, which ensures that the convergence properties of the system in the original space are preserved in the feature space. For the neighbourhood around the equilibrium point $s_{*}$, there exist class $\mathcal{K}$ functions $\alpha_{1}$ and $\alpha_{2}$ as
@@
Thus, when the system is locally stable in the original space, the corresponding system in Hilbert space is also locally stable. According to the Hurwitz stability criterion in A.1, the linear operator $\mathcal{A}$ has only negative real part in its eigenvalues, guaranteeing exponential convergence. If $s_{*}$ is global equilibrium, then $\phi_{S}(s_{*})$ is also a global equilibrium in feature space.
Remark A.1.
The Theorem A.3 (a) implies the existence of a global coordinate. In many situations, when the embedding space is chosen properly, we can have a stronger result that the existence of left inverse such that $\phi^{{\dagger}}\big{(}\phi(s)\big{)}=s$. This result can be naturally connected to the Kazantzis Kravaris/Luenberger (KKL) observers Tran & Bernard (2023). The embedding corresponds the injectivity of the state. The Theorem A.3 (b) corresponds the convergence in KKL observer. When the embedding space is uniformly injective, the dynamics in feature space becomes rectifiable dynamics, yielding a stable trajectory if the original system is stable.
Before entering the last main theorem, we need to introduce the KKL observer. Based on theory of KKL observer, we will link the convergence problem in feature space with KKL observer. Please note, we assume that the state space $\mathcal{S}$ is a bounded set and system equation 1 is forward complete within this bounded set $\mathcal{S}$.
Consider the nonlinear dynamical system in equation 1 with time derivative as
@@
The design of KKL observer is as follows:
  • Find an embedding map $\mathcal{T}:\mathcal{S}\to\mathbb{R}^{d_{s}}$ that transforms equation 17 to new coordinates $\mathcal{T}(s)$ as
@@
There are certain conditions Andrieu & Praly (2006) that equation 17 needs to satisfy in order to ensure the existence of a KKL observer equation 18 in the sense that $\lim_{t\to\infty}\|\hat{s}_{t}-s_{t}\|=0$. On the other hand, the map $\mathcal{T}:S\to\mathbb{R}^{d_{s}}$ need to be uniformly injective if there exists a class $\mathcal{K}$ function $\alpha$ in A.1 such that, for every $s_{1},s_{2}\in S$, satisfying $\|s_{1}-s_{2}\|\leq\alpha(\mathcal{T}(s_{1})-\mathcal{T}(s_{2}))$. Our embedding properties and convergence conditions (as shown in A.3) are satisfied the two conditions, thus we can assert the existence of KKL observer.
In this paper, we state the existence of a global linear dynamical system in feature space. We provide a theoretical guarantee that the embedding property of $\phi_{S}$ can derive the equivalent convergence in the feature space. However, there is two parts that have not been proven: 1). Does the global linear dynamical system exist? 2) Is the embedding space in A.3 properly defined? We consider the nonlinear system in equation 1 within a compact space $S$ and the cost function in 2 has a unique solution. According to the condition of KKL observer, we guarantee (1) the existence of such linear system and (2) the solutions in the original space and feature space has consistent convergence properties, with respect to the cost functions equation 2 and 5, and convergent exponentially to the unique solution.
Theorem A.4.
@@
and where $(s_{t},\mathcal{T}(s_{t}))$ is the solution of system in equation 1 and equation 17 at time $t$; $\sigma_{max}(\mathcal{A})$ is the largest eigenvalue of matrix $\mathcal{A}$.
@@
It yields that the function $\mathcal{T}^{-1}:\mathbb{R}^{d_{s}}\to S$ is global Lipschitz. Hence, the function $\mathcal{T}^{{\dagger}}:\mathbb{R}^{d_{s}}\to S$ to our problem is Lispchitz extension on the set $S$ of this function. For more convenience, we denoted $z\coloneqq\mathcal{T}(s)$. Following the approach - the Mc-Shane formula in McShane (1934), we select the $\mathcal{T}^{{\dagger}}$ as the function defined by
@@
On the another hand, the function $\mathcal{T}$ is solution of the partial differential equation in equation 18, consequently, this implies that along the trajectory of system $(z_{t},s_{t})\in\mathbb{R}^{d_{s}}\times\mathcal{S}$, we have
@@
Note, since $\mathcal{A}$ is Hurwitz matrix, with $\sigma_{max}(\mathcal{A})<0$, it can derive that equation equation 22 holds and concludes the proof of the theorem.
Remark A.2.
The KKL observer asserts that the linear representation of the nonlinear system. After establishing the KKL observer, theorem A.4 asserts the convergence of the estimate trajectory to equilibrium trajectory. Since the embedding property, we can derive that the existence of left inverse based on the Mc-Shane formula in equation 24. Meanwhile, the convergence holds when pulling back the state to original space $\mathcal{S}$. Thus we can assert the feature $\phi_{S}$ in our framework aligns the KKL observer, and the coordinate transformation $\mathcal{T}$ is just our feature $\phi_{S}$.
Appendix B Estimation of error matrices and Pseudo Algorithm
B.1 Estimation of error matrices.
To estimate these error covariance operators in the feature space, we empirically estimate these error matrices in the feature spaces from training dataset (Gejadze et al., 2008). For $\mathcal{R}$ and $\mathcal{Q}$, we estimated the covariances as $\mathcal{R}=\frac{1}{N}\sum_{i=1}^{N}r_{i}\otimes r_{i}$ and $\mathcal{Q}=\frac{1}{N}\sum_{i=1}^{N}q_{i}\otimes q_{i}$, where $r_{i}=\phi_{S}(s_{i})-\hat{\mathcal{C}}_{S|OH}\phi_{OH}(o_{i},h_{i})$ and $q_{i}=\phi_{S}(s^{+}_{i})-\hat{\mathcal{C}}_{S^{+}|S}\phi_{S}(s_{i})$ are the regression residuals, quantifying the errors of the empirical operators $\hat{\mathcal{C}}_{S|OH}$ and $\hat{\mathcal{C}}_{S^{+}|S}$. Similarly, we compute the background covariance as the empirical variance over an average of $\{\phi_{S}(s_{i})\}_{i=1}^{N}$ [^5] For a cyclic application of Tensor-Var, a better design for $\mathcal{B}$ should be time-dependent, reflecting the error between the estimated system state and the true state, e.g. (Paulin et al., 2022). This error should decay monotonically over time and stabilize after a sufficiently long time horizon. This is strongly related to the covariance estimation in Kalman filtering (see Chapter 6.7 (Asch et al., 2016)). We leave the investigation of such design for future efforts.
B.2 Pseudo algorithm.
@@ Appendix B Estimation of error matrices and Pseudo Algorithm
Algorithm 3 outlines procedure of performing data assimilation with trained models.
@@
  Compute the error covariance matrices $\mathcal{B}$, $\mathcal{R}$, $\mathcal{Q}$ from subsection B.1
@@
  Compute the error covariance matrices $\mathcal{B}$, $\mathcal{R}$, $\mathcal{Q}$ from subsection B.1.
@@
Appendix C Experiment Settings
@@
Given the generated data, we constructed two datasets: $\mathcal{D}_{\text{dyn}}=\{\{(s_{t}^{i},s_{t+1}^{i})\}_{t=0}^{T-1}\}_{i=1}^{N}$ and $\mathcal{D}_{\text{obs}}=\{\{(s_{t}^{i},o_{t}^{i},h_{t}^{i})\}_{t=0}^{T}\}_{i=1}^{N}$. The DFs were trained in two steps using these datasets. First, the state DFs $\phi_{\theta_{S}},\phi^{\dagger}_{\theta^{\prime}_{S}}$ were trained on $\mathcal{D}_{\text{dyn}}$ and we stored the estimated operator $\hat{\mathcal{C}}_{S^{+}|S}$. Next, with the state features fixed, the observation DF $\phi_{\theta_{O}}$ and history DF $\phi_{\theta_{H}}$ were trained on $\mathcal{D}_{\text{obs}}$ according to equation 3.2, storing the estimated operator $\hat{\mathcal{C}}_{S|OH}$. The baseline method (Frerix et al., 2021) was trained on $\mathcal{D}_{\text{obs}}$, excluding history. All models were trained with the Adam optimizer (Kingma, 2014) for 200 epochs, using batch sizes from 256 to 1024 for stable operator estimation. Additional details on the DFs, baselines, and training procedures can be found in Appendix C.
@@
For all baseline methods, we employed the L-BFGS algorithm for Variational Data Assimilation (Var-DA) optimization, implemented in JAX (Bradbury et al., 2018). The 4D-Var baselines used numerical dynamical models based on the 8th-order Runge-Kutta method and the 4th-order Exponential Time Differencing Runge-Kutta (ETDRK) method (Cox & Matthews, 2002) for the Lorenz-96 and KS systems. For Tensor-Var, we applied interior-point quadratic programming to solve the linearized 4D-Var optimization, utilizing CVXPY (Diamond & Boyd, 2016). All training was conducted on a workstation with a 48-core AMD 7980X CPU and an Nvidia GeForce 4090 GPU.
C.1 Lorenz 96
First, we consider the single-level Lorenz-96 system, which was introduced in (Lorenz, 1996) as a low-order model of atmospheric circulation along a latitude circle. The system state is $[S_{1},...,S_{K}]$ representing atmospheric velocity at $K$ evenly spaced locations and is evolved according to the governing equation:

### v3

@@ Appendix C Theoretical Analysis
In this section, we provide a theoretical convergence analysis of Tensor-Var, drawing on concepts from control theory and contraction analysis. We begin by introducing comparison functions, including class $\mathcal{K}$ and $\mathcal{KL}$ functions, as well as Lyapunov functions. Using these tools, we examine the convergence of Tensor-Var through a differential equation, demonstrating monotonic contraction based on the Lyapunov direct method. Furthermore, by using comparison functions, we show that contraction in the feature space implies contraction in the original space.
@@
Assumption 1. To perform a formal convergence analysis of Tensor-Var, we make a mild assumption about the first-order differentiability of the dynamical system $F$ in (1) with the time derivative $\dot{s}=f(s)$, a standard assumption in convergence studies (Sastry, 2013).
@@
C.1 Notations and Technical Lemmas
Definition C.1 (Class $\mathcal{K}$ function (Gajic & Qureshi, 2008).).
@@ C.1 Notations and Technical Lemmas
Definition C.2 (Class $\mathcal{KL}$ function (Gajic & Qureshi, 2008).).
A continuous function $\beta:[0,a]\times[0,\infty)\to[0,\infty)$ is said to belong to class $\mathcal{KL}$ if for each fixed $t$, the mapping $\beta(r,t)$ belonging to class $\mathcal{K}$ with respect to $r$ and, for each fixed $a$ the mapping $\beta(a,t)$ is decreasing with respect to $t$ and $\beta(a,t)\to 0$ as $t\to\infty$.
The $\mathcal{K}$ and $\mathcal{KL}$ are two classes of comparison functions, and we can use the comparison function to analyze the monotone contraction in both spatial and temporal horizons.
Definition C.3 (Lyapunov Stability (Gajic & Qureshi, 2008)).
If the Lyapunov function $V$ is globally positive definite, radially unbounded, the equilibrium is isolated and the time derivative of the Lyapunov function is globally negative definite:
@@
then the equilibrium is proven to be globally asymptotically stable. The Lyapunov function is a class $\mathcal{K}$ function, which satisfies the condition as follows
@@
Lemma C.4 (Hurwitz stability criterion (Parks, 1962)).
@@
Lemma C.5 (Hurwitz stability criterion via Lyapunov function (Sastry, 2013)).
@@
Definition C.6 (Contraction (Manchester & Slotine, 2017)).
@@
C.2 Proof of Key Theorems
To adopt the convergence analysis of the DA problem, we give a mild assumption on the smoothness of dynamics, the first-order derivative exists in the system 1.
Theorem C.7 (Embedding and Consistent convergence).
@@ C.2 Proof of Key Theorems
The second line of (LABEL:Equation:_dynamical_system_in_Hilbert_space) follows from the chain rule, the final line represents the time derivative in Hilbert space, where $\mathcal{A}$ is a linear operator that governs the dynamics. Following the work from (Romanoff, 1947; Bobrowski, 2016; Cheng et al., 2023b), we define $\mathcal{A}$ as:
@@
where $\mathcal{C}_{S^{+}\mid S}$ is the conditional covariance operator between future and current states in the RKHS. Given the smoothness of the kernel function and the differentiability of the system dynamics, the linear operator $\mathcal{A}$ exists and is well-defined in this context.
Since $s_{*}$ is the equilibrium point, we can derive a natural result that $f(s_{*})=0$. Invoking (LABEL:Equation:_dynamical_system_in_Hilbert_space) (third line) we have $\frac{d\phi_{S}}{dt}(s_{*})=0$. Thus $\phi_{S}(s_{*})$ is also a local equilibrium point in RKHS. From the embedding property of RKHS, we have a local injective map $\phi:\mathbb{R}^{n_{s}}\to\mathbb{R}^{d_{s}}$, which ensures that the convergence properties of the system in the original space are preserved in the feature space. For the neighbourhood around the equilibrium point $s_{*}$, there exist class $\mathcal{K}$ functions $\alpha_{1}$ and $\alpha_{2}$ as
@@
Thus, when the system is locally stable in the original space, the corresponding system in Hilbert space is also locally stable. According to the Hurwitz stability criterion in C.4, the linear operator $\mathcal{A}$ has only negative real part in its eigenvalues, guaranteeing exponential convergence. If $s_{*}$ is global equilibrium, then $\phi_{S}(s_{*})$ is also a global equilibrium in feature space.
Remark C.8.
Theorem C.7 (a) implies the existence of a global coordinate. In many situations, when the embedding space is chosen properly, we can have a stronger result that the existence of left inverse such that $\phi^{{\dagger}}\big{(}\phi(s)\big{)}=s$. This result can be naturally connected to the Kazantzis Kravaris/Luenberger (KKL) observers (Tran & Bernard, 2023). The embedding corresponds to the injectivity of the state. Theorem C.7 (b) corresponds to the convergence in the KKL observer. When the embedding space is uniformly injective, the dynamics in feature space become rectifiable dynamics, yielding a stable trajectory if the original system is stable.
Before entering the last main theorem, we need to introduce the KKL observer. Based on the theory of the KKL observer, we will link the convergence problem in feature space with the KKL observer. Please note that we assume that the state space $\mathcal{S}$ is a bounded set and system (1) is forward complete within this bounded set $\mathcal{S}$.
Consider the nonlinear dynamical system in (1) with time derivative as
@@
The design of the KKL observer is as follows:
  • Find an embedding map $\mathcal{T}:\mathcal{S}\to\mathbb{R}^{d_{s}}$ that transforms (17) to new coordinates $\mathcal{T}(s)$ as
@@
There are certain conditions (Andrieu & Praly, 2006) that (17) needs to satisfy in order to ensure the existence of a KKL observer (18) in the sense that $\lim_{t\to\infty}\|\hat{s}_{t}-s_{t}\|=0$. On the other hand, the map $\mathcal{T}:S\to\mathbb{R}^{d_{s}}$ need to be uniformly injective if there exists a class $\mathcal{K}$ function $\alpha$ in C.1 such that, for every $s_{1},s_{2}\in S$, satisfying $\|s_{1}-s_{2}\|\leq\alpha(\mathcal{T}(s_{1})-\mathcal{T}(s_{2}))$. Our embedding properties and convergence conditions (as shown in C.7) are satisfied by the two conditions, thus we can assert the existence of a KKL observer.
In this paper, we state the existence of a global linear dynamical system in feature space. We provide a theoretical guarantee that the embedding property of $\phi_{S}$ can derive the equivalent convergence in the feature space. However, there are two parts that have not been proven: 1). Does the global linear dynamical system exist? 2) Is the embedding space in C.7 properly defined? We consider the nonlinear system in (1) within a compact space $S$ and the cost function in 2 has a unique solution. According to the condition of KKL observer, we guarantee (1) the existence of such linear system and (2) the solutions in the original space and feature space has consistent convergence properties, with respect to the cost functions (2) and 5, and convergent exponentially to the unique solution.
Theorem C.9.
@@
and where $(s_{t},\mathcal{T}(s_{t}))$ is the solution of system in (1) and (17) at time $t$; $\sigma_{max}(\mathcal{A})$ is the largest eigenvalue of matrix $\mathcal{A}$.
@@
It yields that the function $\mathcal{T}^{-1}:\mathbb{R}^{d_{s}}\to S$ is global Lipschitz. Hence, the function $\mathcal{T}^{{\dagger}}:\mathbb{R}^{d_{s}}\to S$ in our problem is a Lipschitz extension on the set $S$ of this function. For more convenience, we denoted $z\coloneqq\mathcal{T}(s)$. Following the approach - the Mc-Shane formula in (McShane, 1934), we select the $\mathcal{T}^{{\dagger}}$ as the function defined by
@@
On the another hand, the function $\mathcal{T}$ is solution of the partial differential equation in (18), consequently, this implies that along the trajectory of system $(z_{t},s_{t})\in\mathbb{R}^{d_{s}}\times\mathcal{S}$, we have
@@
Note, since $\mathcal{A}$ is Hurwitz matrix, with $\sigma_{max}(\mathcal{A})<0$, it can derive that equation (22) holds and concludes the proof of the theorem.
Remark C.10.
The KKL observer asserts that the linear representation of the nonlinear system. After establishing the KKL observer, theorem C.9 asserts the convergence of the estimated trajectory to the equilibrium trajectory. Since the embedding property, we can derive that the existence of left inverse based on the Mc-Shane formula in (24). Meanwhile, the convergence holds when pulling back the state to the original space $\mathcal{S}$. Thus we can assert the feature $\phi_{S}$ in our framework aligns the KKL observer, and the coordinate transformation $\mathcal{T}$ is just our feature $\phi_{S}$.
Appendix D Estimation of error matrices and Pseudo Algorithm
D.1 Estimation of error matrices.
To estimate these error covariance operators in the feature space, we empirically estimate these error matrices in the feature spaces from the training dataset (Gejadze et al., 2008). For $\mathcal{R}$ and $\mathcal{Q}$, we estimated the covariances as $\mathcal{R}=\frac{1}{N}\sum_{i=1}^{N}r_{i}\otimes r_{i}$ and $\mathcal{Q}=\frac{1}{N}\sum_{i=1}^{N}q_{i}\otimes q_{i}$, where $r_{i}=\phi_{S}(s_{i})-\hat{\mathcal{C}}_{S|OH}\phi_{OH}(o_{i},h_{i})$ and $q_{i}=\phi_{S}(s^{+}_{i})-\hat{\mathcal{C}}_{S^{+}|S}\phi_{S}(s_{i})$ are the regression residuals, quantifying the errors of the empirical operators $\hat{\mathcal{C}}_{S|OH}$ and $\hat{\mathcal{C}}_{S^{+}|S}$. Similarly, we compute the background covariance as the empirical variance over an average of $\{\phi_{S}(s_{i})\}_{i=1}^{N}$ [^5] For a cyclic application of Tensor-Var, a better design for $\mathcal{B}$ should be time-dependent, reflecting the error between the estimated system state and the true state, e.g. (Paulin et al., 2022). This error should decay monotonically over time and stabilize after a sufficiently long time horizon. This is strongly related to the covariance estimation in Kalman filtering (see Chapter 6.7 (Asch et al., 2016)). We leave the investigation of such design for future efforts.
D.2 Pseudo algorithm.
@@ Appendix D Estimation of error matrices and Pseudo Algorithm
Algorithm 3 outlines the procedure of performing data assimilation with trained models.
@@
  Compute the error covariance matrices $\mathcal{B}$, $\mathcal{R}$, $\mathcal{Q}$ from Subsection D.1
@@
  Compute the error covariance matrices $\mathcal{B}$, $\mathcal{R}$, $\mathcal{Q}$ from Subsection D.1.
@@
Appendix E Experiment Settings
@@
Given the generated data, we construct two datasets: $\mathcal{D}_{\text{dyn}}=\{\{(s_{t}^{i},s_{t+1}^{i})\}_{t=0}^{T-1}\}_{i=1}^{N}$ and $\mathcal{D}_{\text{obs}}=\{\{(s_{t}^{i},o_{t}^{i},h_{t}^{i})\}_{t=0}^{T}\}_{i=1}^{N}$. The DFs are trained in two steps using these datasets. First, the state DFs $\phi_{\theta_{S}},\phi^{\dagger}_{\theta^{\prime}_{S}}$are trained on $\mathcal{D}_{\text{dyn}}$ and we store the estimated operator $\hat{\mathcal{C}}_{S^{+}|S}$. Next, with the state features fixed, the observation DF $\phi_{\theta_{O}}$ and history DF $\phi_{\theta_{H}}$ are trained on $\mathcal{D}_{\text{obs}}$ according to (3.2), storing the estimated operator $\hat{\mathcal{C}}_{S|OH}$. The baseline method (Frerix et al., 2021) is trained on $\mathcal{D}_{\text{obs}}$, excluding history. All models are trained with the Adam optimizer (Kingma, 2014) for 200 epochs, using batch sizes from 256 to 1024 for stable operator estimation. Additional details on the DFs, baselines, and training procedures can be found in Appendix E.
@@
For all baseline methods, we employ the L-BFGS algorithm for Variational Data Assimilation (Var-DA) optimization, implemented in JAX (Bradbury et al., 2018). The 4D-Var baselines use numerical dynamical models based on the 8th-order Runge-Kutta method and the 4th-order Exponential Time Differencing Runge-Kutta (ETDRK) method (Cox & Matthews, 2002) for the Lorenz-96 and KS systems. For Tensor-Var, we apply interior-point quadratic programming to solve the linearized 4D-Var optimization, using CVXPY (Diamond & Boyd, 2016). All training is conducted on a workstation with a 48-core AMD 7980X CPU and an Nvidia GeForce 4090 GPU.
E.1 Lorenz 96
First, we consider the single-level Lorenz-96 system, which is introduced in (Lorenz, 1996) as a low-order model of atmospheric circulation along a latitude circle. The system state is $[S_{1},...,S_{K}]$ representing atmospheric velocity at $K$ evenly spaced locations and is evolved according to the governing equation: